# Does Graph Prompt Work? A Data Operation Perspective with Theoretical Analysis

### v1

@@ A.2.3 On the Range of Graph Embedding Matrix After Graph Prompting, Not Full-Rank Case
At this point, we can consider $\hat{R}^{\prime}_{1}$ equal to zero. Since, for GPF, $\mathbf{c}_{0}=\mathbf{1}_{N}$, and the monotonicity of $\mathbf{c}$ implies that each component of the vector $\mathbf{c}$ is strictly greater than 0. For All-in-One, $\mathbf{c}=\mathbf{e}_{N}$. We have $S_{\omega}=[s_{1},\ldots,s_{N}]$, where $s_{N}=\begin{pmatrix}l\\ S_{NN}\end{pmatrix}$, $\mathbf{c}^{\prime}=s_{N}$. we assume $l>0$ component-wise(This can be achieved by adjusting parameters in All in one), then $\mathbf{c}^{\prime}$ is component-wise $>0$.
@@
$\sigma(S_{\omega}(\hat{R}^{\prime}_{1}+\mathbf{c}_{1}\mathbf{p}^{\top})W)=\sigma(S_{\omega}(\mathbf{c}_{1}\mathbf{p}^{\top})W)=\sigma(\mathbf{c}_{2}\mathbf{p^{\prime}}^{\top})=\mathbf{c}_{2}\mathbf{p^{\prime\prime}}^{\top}$ (36)
@@ A.3.3 On Error Upper Bound Analysis of Mapping to Bridge Set at Single Graph Level
$\|F_{\theta}(P_{\omega}(G_{ori}))-C(G_{ori})\|$ $=\|F_{\theta}(P_{\omega}(G_{ori}))-F_{\theta}(G_{bri})\|$ (40)
@@ A.3.4 Error Bound Analysis of Mapping to Bridge Set at Multiple Graph Level
$F_{\theta}(G_{i})=\mathbf{w}^{\top}H_{i}^{(n)}=\mathbf{w}^{\top}R_{i}^{(n)}+\mathbf{w}^{\top}\mathbf{c}_{i}\mathbf{p}^{\top}=R_{0}^{\top}+\lambda_{i}\mathbf{p}^{\top}$ (44)
@@
The optimal $p$ is the weighted centroid of these $n$ vectors $(C_{\lambda_{i}}(G_{i}))_{i=1}^{M}$, as pointed out at Boyd & Vandenberghe (2004), i.e.:
@@
$I(G_{1},\ldots,G_{M})$ $=\min_{\omega}(\sqrt{\sum_{i=1}^{M}\lambda_{i}^{2}\|C_{\lambda_{i}}(G_{i})-\mathbf{p}\|^{2}/M})=\sqrt{\sum_{i=1}^{M}\|C(G_{i})-\lambda_{i}\mathbf{p}^{*}\|^{2}/M}$ (50)
@@ A.3.5 Error Distribution Analysis of Mapping to Bridge Set
Consider $K_{\theta}(G_{bri})=H^{(n)}(G_{bri})=\begin{pmatrix}\mathbf{v}^{\top}_{1}\\ \vdots\\ \mathbf{v}^{\top}_{n}\end{pmatrix}$, where $\mathbf{v}_{i}\in\mathbb{R}^{F}$.
Then, $\Delta F_{\theta}(G_{ori})=F_{\theta}(G_{\omega})-C(G_{ori})=\mathbf{w}^{\top}H^{(n)}(G_{\omega})-\mathbf{w}^{\top}H^{(n)}(G_{bri})=\mathbf{w}^{\top}(H^{(n)}(G_{\omega})-H^{(n)}(G_{bri}))=\mathbf{w}^{\top}\Delta H$.
@@
Let $p^{\prime*}=\begin{pmatrix}\alpha_{1}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&\alpha_{F}\end{pmatrix}p^{*}$ and $\bm{\beta}=[\beta_{1},\ldots,\beta_{F}]^{\top}$.
Then $p^{\prime*}\in R(W)$, so $\|\Delta F_{\theta}(G)\|=\|p^{\prime*}-\bm{\beta}\|$. This value represents the distance between $\bm{\beta}$ and $p^{\prime*}$. Since $p^{\prime*}$ already minimizes $\|\bm{A}p^{\prime}-\bm{\beta}\|$ and $\bm{A}p^{\prime}\in R(\bm{A}W)$, $\forall p^{\prime}\in R(W)$, where $\bm{A}$ denotes $diag(\alpha_{1},\cdots,\alpha_{F})$,we can estimate $\|p^{*^{\prime}}-\bm{\beta}\|$ using the distance from $\bm{\beta}$ to the space $R(\bm{A}W)$.
Consider the projection matrix $P$ onto the space $R(\bm{A}W)$. Then $\|\Delta F_{\theta}(G)\|=\|(I-P)\bm{\beta}\|$.
@@
$\beta_{j}=\sum^{N}_{i=1}-(\alpha)^{k}w_{i}v_{ij},\quad\text{where }k=\begin{cases}0&\text{if }\sigma(c_{i}p^{\prime}_{j}-\sigma^{-1}(v_{ij}))>0\\ -1&\text{otherwise}\end{cases}$ (60)
Where $\alpha$ is the is the parameter of the Leaky-ReLU. According to the Central Limit Theorem and the independence of different components in $\bm{\beta}$, it is reasonable to assume that $\bm{\beta}$ follows an $n$-dimensional i.i.d. normal distribution with mean 0 and variance $c$ for some positive constant $c$ as discussed in 5.
@@
Here, our equation 58 implicitly assumes that the first $N-1$ layers of GNN are of full rank. However, even if the rank of the first $N-1$ layers of GNN could be non-full-rank, we still have $\mathbf{p^{\prime}}\in R(W)$, according to lemma 3. Therefore, we can still use this theorem for approximation.
@@
We assume that the vector $\bm{\beta}$, which we call the graph embedding residual vector, follows an $F$-dimensional i.i.d. normal distribution with mean $\mathbf{0}$ and variance $c$:
$\bm{\beta}\sim\mathcal{N}(0,cI_{F})$ (61)
@@
Intuitively, this graph embedding residual vector $\bm{\beta}$ represents the additional term that the graph needs to fit due to non-full-rank and non-linear components. Formally, its components are defined as:
$[\bm{\beta}]_{j}=\sum^{N}_{i=1}-(\alpha)^{k}w_{i}v_{ij},\quad\text{where }k=\begin{cases}0&\text{if }\sigma(c_{i}p^{\prime}_{j}-\sigma^{-1}(v_{ij}))>0\\ -1&\text{otherwise}\end{cases}$ (62)
The reasons behind this assumption are as follows: (1). Symmetry (2). The sum of random variables and Central Limit Theorem: Each $\beta_{j}$ is a sum of multiple terms, each of which can be considered a random variable. The Central Limit Theorem suggests that its distribution should approach a normal distribution. (3). Continuous and smooth distribution (4). Independence: The components of $\bm{\beta}$ are assumed to be independent due to the independence of the input features and the structure of the GNN.
Therefore, it is reasonable to assume that $\bm{\beta}$ follows an $F$-dimensional i.i.d. normal distribution with mean 0 and variance $c$ for some positive constant $c$.
@@ A.3.6 Analysis of Nonlinear Graph Neural Networks
$H^{(m)}=\sigma((S_{\omega}\odot\langle X_{\omega},X_{\omega}\rangle)\cdot X_{\omega}\cdot W)$
@@
$S_{\omega}\odot\langle X_{\omega},X_{\omega}\rangle W$ $=\begin{pmatrix}S\odot\langle X,X\rangle&l\odot(X\cdot\mathbf{p})\\ l^{T}\odot(X\cdot\mathbf{p})^{T}&0\end{pmatrix}\cdot\begin{pmatrix}X\\ \mathbf{p}^{\top}\end{pmatrix}\cdot W$ (63)
@@
$H[1:N,]$ $=[\begin{pmatrix}S\odot\langle X,X\rangle&X\cdot(l\odot\mathbf{p})\\ l^{T}\odot(X\cdot\mathbf{p})^{T}&(\mathbf{p}^{\top}\mathbf{p})\end{pmatrix}\cdot\begin{pmatrix}X\\ \mathbf{p}^{\top}\end{pmatrix}\cdot W][1:N,]$ (65)
@@ A.4 Further Mathematical Discussion
$p_{j}^{\prime}(p_{j})=\sum^{N}_{i=1}w_{i}(R_{ij}(p_{j})+c_{ij}(p_{j})p_{j})=R^{\prime}_{j}(p_{j})+C^{\prime}_{j}(p_{j})p_{j}$
@@
$[C(\mathbf{p_{i}})]_{j}=[\text{ReadOut}(\sigma(\mathbf{cp_{i}}^{\top}))]_{j}=\sum^{N}_{i=1}w_{i}c_{i}\sigma([\mathbf{p_{i}}]_{j})=\lambda\sigma([\mathbf{p_{i}}]_{j})$
@@
$\tilde{C}$ $=\text{diag}(\lambda_{1},\ldots,\lambda_{n})\begin{pmatrix}[\mathbf{p_{1}}]_{1}\cdots[\mathbf{p_{1}}]_{F}\\ \vdots\\ [\mathbf{p_{M}}]_{1}\cdots[\mathbf{p_{M}}]_{F}\\ \end{pmatrix}\text{diag}(\alpha^{I_{1}},\cdots,\alpha^{I_{F}})$ (77)
@@
$\quad-VV[:,:k]V[:,:k]^{T}\Lambda V^{T}+VV[:,:k]V[:,:k]^{T}\Lambda V^{T}V[:,:k]V[:,:k]^{T})$
@@ Appendix B Supplementary Experiments
We conducted supplementary experiments on real-world datasets to validate our results. We performed similar experiments as in the main text using entirely different datasets, namely NCI1 and DD Morris et al. (2020). The NCI1 dataset has 37-dimensional node features, 4,110 subgraphs, an average of 29.87 nodes per subgraph, and an average of 64.60 edges per subgraph. The DD dataset has 89-dimensional node features, 1,178 subgraphs, an average of 284.32 nodes per subgraph, and an average of 1,431.32 edges per subgraph. Our results are shown in the following figures.
@@
Figure 6: convergence analysis
@@
Figure 7: Empirical Max Error v.s. Rank Lost
@@
(a) NCI1
(b) DD
Figure 8: Empirical Min Error v.s. Graph Number
@@ Appendix C EXPERIMENT DETAILS
Model Settings: We utilize two GNN frameworks: GCN (representing linear models) and GAT (representing non-linear models). We limit our experiments to these two models as other models follow similar patterns. Unless otherwise specified, we use a 3-layer GNN with Leaky-ReLU activation function and feature dimension $F=25$. For full-rank matrix studies, we ensure each layer’s matrix is full-rank (selected after pre-training). For non-full-rank matrix studies, we set the rank loss to $5$ by default. The default ReadOut method is mean pooling.

### v2

@@ A.2.3 On the Range of Graph Embedding Matrix After Graph Prompting, Not Full-Rank Case
At this point, we can consider $\hat{R}^{\prime}_{1}$ equal to zero. Since, for GPF, $\mathbf{c}_{0}=\mathbb{1}_{N}$, and the monotonicity of $\mathbf{c}$ implies that each component of the vector $\mathbf{c}$ is strictly greater than 0. For All-in-One, $\mathbf{c}=\hat{\mathbf{e}}_{N}$. We have $S_{\omega}=[s_{1},\ldots,s_{N}]$, where $s_{N}=\begin{pmatrix}l\\ S_{NN}\end{pmatrix}$, $\mathbf{c}^{\prime}=s_{N}$. we assume $l>0$ component-wise(This can be achieved by adjusting parameters in All in one), then $\mathbf{c}^{\prime}$ is component-wise $>0$.
@@
$\sigma(S_{\omega}(\hat{R}^{\prime}_{1}+\mathbf{c}_{1}\mathbf{p}^{\top})W)$ $=\sigma(S_{\omega}(\mathbf{c}_{1}\mathbf{p}^{\top})W)$ (36)
$=\sigma(\mathbf{c}_{2}\mathbf{p^{\prime}}^{\top})$ $=\mathbf{c}_{2}\mathbf{p^{\prime\prime}}^{\top}$ (37)
@@ A.3.3 On Error Upper Bound Analysis of Mapping to Bridge Set at Single Graph Level
$\|F_{\theta}(P_{\omega}(G_{ori}))-C(G_{ori})\|$ (41)
$=\|F_{\theta}(P_{\omega}(G_{ori}))-F_{\theta}(G_{bri})\|$ (42)
@@ A.3.4 Error Bound Analysis of Mapping to Bridge Set at Multiple Graph Level
$F_{\theta}(G_{i})=\mathbf{w}^{\top}H_{i}^{(n)}$ (46)
$=\mathbf{w}^{\top}R_{i}^{(n)}+\mathbf{w}^{\top}\mathbf{c}_{i}\mathbf{p}^{\top}=R_{0}^{\top}+\lambda_{i}\mathbf{p}^{\top}$ (47)
@@
The optimal $p$ is the weighted centroid of these $n$ vectors $(C_{\lambda_{i}}(G_{i}))_{i=1}^{M}$, as pointed out at (Boyd & Vandenberghe, 2004), i.e.:
@@
$I(G_{1},\ldots,G_{M})$ (53)
$=\min_{\omega}(\sqrt{\sum_{i=1}^{M}\lambda_{i}^{2}\|C_{\lambda_{i}}(G_{i})-\mathbf{p}\|^{2}/M})$ (54)
$=\sqrt{\sum_{i=1}^{M}\|C(G_{i})-\lambda_{i}\mathbf{p}^{*}\|^{2}/M}$ (55)
@@ A.3.5 Error Distribution Analysis of Mapping to Bridge Set
Consider $K_{\theta}(G_{bri})=H^{(n)}(G_{bri})=\begin{pmatrix}\mathbf{v}^{\top}_{1}\\ \vdots\\ \mathbf{v}^{\top}_{n}\end{pmatrix}$, where $\mathbf{v}_{i}\in\mathbb{R}^{F}$. Then,
$\Delta F_{\theta}(G_{ori})$ $=F_{\theta}(G_{\omega})-C(G_{ori})$
$=\mathbf{w}^{\top}H^{(n)}(G_{\omega})-\mathbf{w}^{\top}H^{(n)}(G_{bri})$
$=\mathbf{w}^{\top}\big{(}H^{(n)}(G_{\omega})-H^{(n)}(G_{bri})\big{)}$
$=\mathbf{w}^{\top}\Delta H.$ (65)
@@
Let $p^{\prime*}=\begin{pmatrix}\alpha_{1}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&\alpha_{F}\end{pmatrix}p^{*}$ and $\boldsymbol{\beta}=[\beta_{1},\ldots,\beta_{F}]^{\top}$.
Then $p^{\prime*}\in R(W)$, hence $\|\Delta F_{\theta}(G)\|=\|p^{\prime*}-\boldsymbol{\beta}\|$. This value represents the distance between $\boldsymbol{\beta}$ and $p^{\prime*}$. Since $p^{\prime*}$ already minimizes $\|\boldsymbol{A}p^{\prime}-\boldsymbol{\beta}\|$ and $\boldsymbol{A}p^{\prime}\in R(\boldsymbol{A}W)$, $\forall p^{\prime}\in R(W)$, where $\boldsymbol{A}$ denotes $diag(\alpha_{1},\cdots,\alpha_{F})$,we can estimate $\|p^{*^{\prime}}-\boldsymbol{\beta}\|$ using the distance from $\boldsymbol{\beta}$ to the space $R(\boldsymbol{A}W)$.
Consider the projection matrix $P$ onto the space $R(\boldsymbol{A}W)$. Then $\|\Delta F_{\theta}(G)\|=\|(I-P)\boldsymbol{\beta}\|$.
@@
$\beta_{j}=\sum^{N}_{i=1}-(\alpha)^{k}w_{i}v_{ij}$ (66)
Where, $k=\begin{cases}0&\text{if }\sigma(c_{i}p^{\prime}_{j}-\sigma^{-1}(v_{ij}))>0\\ -1&\text{otherwise}\end{cases}$ (67)
Where $\alpha$ is the is the parameter of the Leaky-ReLU. According to the Central Limit Theorem and the independence of different components in $\boldsymbol{\beta}$, it is reasonable to assume that $\boldsymbol{\beta}$ follows an $n$-dimensional i.i.d. normal distribution with mean 0 and variance $c$ for some positive constant $c$ as discussed in 5.
@@
Here, our (63) implicitly assumes that the first $N-1$ layers of GNN are of full rank. However, even if the rank of the first $N-1$ layers of GNN could be non-full-rank, we still have $\mathbf{p^{\prime}}\in R(W)$, according to lemma 3. Therefore, we can still use this theorem for approximation.
[…]
@@
We assume that the vector $\boldsymbol{\beta}$, which we call the graph embedding residual vector, follows an $F$-dimensional i.i.d. normal distribution with mean $\mathbf{0}$ and variance $c$:
$\boldsymbol{\beta}\sim\mathcal{N}(0,cI_{F})$ (68)
@@
Intuitively, this graph embedding residual vector $\boldsymbol{\beta}$ represents the additional term that the graph needs to fit due to non-full-rank and non-linear components. Formally, its components are defined as:
$[\boldsymbol{\beta}]_{j}=\sum^{N}_{i=1}-(\alpha)^{k}w_{i}v_{ij}$ (69)
Where, $k=\begin{cases}0&\text{if }\sigma(c_{i}p^{\prime}_{j}-\sigma^{-1}(v_{ij}))>0\\ -1&\text{otherwise}\end{cases}$ (70)
The reasons behind this assumption are as follows: (1). Symmetry (2). The sum of random variables and Central Limit Theorem: Each $\beta_{j}$ is a sum of multiple terms, each of which can be considered a random variable. The Central Limit Theorem suggests that its distribution should approach a normal distribution. (3). Continuous and smooth distribution (4). Independence: The components of $\boldsymbol{\beta}$ are assumed to be independent due to the independence of the input features and the structure of the GNN.
Therefore, it is reasonable to assume that $\boldsymbol{\beta}$ follows an $F$-dimensional i.i.d. normal distribution with mean 0 and variance $c$ for some positive constant $c$.
@@ A.3.6 Analysis of Nonlinear Graph Neural Networks
$H=\sigma((S_{\omega}\odot\langle X_{\omega},X_{\omega}\rangle)\cdot X_{\omega}\cdot W)$
@@
$S_{\omega}\odot\langle X_{\omega},X_{\omega}\rangle W$ (71)
$=\begin{pmatrix}S\odot\langle X,X\rangle&l\odot(X\cdot\mathbf{p})\\ l^{T}\odot(X\cdot\mathbf{p})^{T}&0\end{pmatrix}\cdot\begin{pmatrix}X\\ \mathbf{p}^{\top}\end{pmatrix}\cdot W$ (72)
@@
$H[1:N,]$ (74)
$=[\begin{pmatrix}S\odot\langle X,X\rangle&X\cdot(l\odot\mathbf{p})\\ l^{T}\odot(X\cdot\mathbf{p})^{T}&(\mathbf{p}^{\top}\mathbf{p})\end{pmatrix}\cdot\begin{pmatrix}X\\ \mathbf{p}^{\top}\end{pmatrix}\cdot W][1:N,]$ (75)
@@ A.4 Further Mathematical Discussion
$p_{j}^{\prime}(p_{j})$ $=\sum^{N}_{i=1}w_{i}(R_{ij}(p_{j})+c_{ij}(p_{j})p_{j})$
$=R^{\prime}_{j}(p_{j})+C^{\prime}_{j}(p_{j})p_{j}$
@@
$[C(\mathbf{p_{i}})]_{j}$ $=[\text{ReadOut}(\sigma(\mathbf{cp_{i}}^{\top}))]_{j}$
$=\sum^{N}_{i=1}w_{i}c_{i}\sigma([\mathbf{p_{i}}]_{j})=\lambda\sigma([\mathbf{p_{i}}]_{j})$
@@
$\tilde{C}=\begin{pmatrix}\lambda_{1}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&\lambda_{n}\end{pmatrix}\begin{pmatrix}[\mathbf{p_{1}}]_{1}&\cdots&[\mathbf{p_{1}}]_{F}\\ \vdots&\ddots&\vdots\\ [\mathbf{p_{M}}]_{1}&\cdots&[\mathbf{p_{M}}]_{F}\\ \end{pmatrix}\begin{pmatrix}\alpha^{I_{1}}&\cdots&0\\ \vdots&\ddots&\vdots\\ 0&\cdots&\alpha^{I_{F}}\end{pmatrix}$ (87)
@@
$\quad-VV[:,:k]V[:,:k]^{T}\Lambda V^{T}+$
$VV[:,:k]V[:,:k]^{T}\Lambda V^{T}V[:,:k]V[:,:k]^{T})$
@@ Appendix B Supplementary Experiments
We conducted supplementary experiments on real-world datasets to validate our results. We performed similar experiments as in the main text using entirely different datasets, namely NCI1 and DD (Morris et al., 2020). The NCI1 dataset has 37-dimensional node features, 4,110 subgraphs, an average of 29.87 nodes per subgraph, and an average of 64.60 edges per subgraph. The DD dataset has 89-dimensional node features, 1,178 subgraphs, an average of 284.32 nodes per subgraph, and an average of 1,431.32 edges per subgraph. Our results are shown in the following figures.
@@
Figure 6: Convergence analysis.
@@
Figure 7: Empirical Max Error v.s. Rank Lost.
@@
(a) NCI1                          (b) DD
Figure 8: Empirical Min Error v.s. Graph Number.
@@
Figure 9: Additional multiple graph analysis with All-in-one-plus
Exp 4 we implemented Experiment B in Section 5.4 to supplement the analysis of All-in-one-plus in multi-graph prompts. All-in-one-plus also achieved a reduction in overall error due to the expanded search space, and the resulting surface exhibits a shape similar to that of the other two experiments in Section 5.4(b).
@@ Appendix C EXPERIMENT DETAILS
Model Settings: To evaluate our approach, we conduct experiments using two representative Graph Neural Network (GNN) architectures: Graph Convolutional Networks (GCN) as linear propagation models and Graph Attention Networks (GAT) as non-linear propagation models. We focus on these two architectures because other models tend to exhibit similar behavior patterns. Unless stated otherwise, our default configuration employs a three-layer GNN with a feature dimension of $F=25$ and utilizes the Leaky-ReLU activation function. For experiments involving full-rank matrices, we ensure that each layer’s weight matrix is full rank, selected after an initial pre-training phase. Conversely, for studies with non-full-rank matrices, we set the rank loss parameter to a default value of $5$. We adopt mean pooling as the default readout method throughout our experiments.
@@
Definition of $C(G)$. In our experiments, $C(G)$ represents the optimal graph-level embedding of a modified graph $G^{\prime}$ derived from the original graph $G$. Specifically, $G^{\prime}$ is obtained by performing a graph data manipulation operation on $G$, such as removing a certain percentage $\epsilon\%$ of edges or nodes. The function $C(G)$ is defined as:$C(G)=\text{Pooling}(\text{GNN}(G^{\prime}))$ where GNN is a graph neural network, and Pooling is a graph-level pooling operation that aggregates node embeddings into a single vector representing the entire graph $G^{\prime}$. This embedding $C(G)$ serves as the ground truth in our experiments.
Experimental Procedure. We focus on a graph-level task that inherently requires graph data manipulation. The experimental procedure is as follows:
  1. 1.
    Graph Manipulation: Starting with an original graph $G$, we generate a modified graph $G^{\prime}$ by randomly applying graph manipulations, including all the aforementioned actions (specifically, adding/deleting nodes, adding/deleting/changing edges, and transforming features of $G$, each with chosen hyperparameter ratios). This process simulates a data operation that alters the graph structure.
  2. 2.
    Computing Ground Truth Embedding $C(G)$: We pass $G^{\prime}$ through a pre-trained and fixed GNN model followed by a pooling operation to obtain the graph-level embedding $C(G)$: $G^{\prime}\rightarrow\text{GNN}\rightarrow\text{Pooling}\rightarrow C(G)$ This embedding represents the desired outcome of the data operation.
  3. 3.
    Graph Prompting on $G$: Instead of directly manipulating $G$, we apply a graph prompt $P_{w}$ to the original graph $G$ to approximate the effect of the manipulation: $G\rightarrow P_{w}(G)$ Here, $P_{w}(G)$ is the prompted graph, where $P_{w}$ is a learnable function parameterized by $w$.
  4. 4.
    Computing Prompted Graph Embedding: We pass the prompted graph $P_{w}(G)$ through the same GNN and pooling operations to obtain the embedding $F_{\theta}(P_{w}(G))$: $P_{w}(G)\rightarrow\text{GNN}\rightarrow\text{Pooling}\rightarrow F_{\theta}(P_{w}(G))$
  5. 5.
    Error Computation: We compute the error between the embeddings of the prompted graph and the ground truth embedding: $\text{Error}=\left\|F_{\theta}(P_{w}(G))-C(G)\right\|$ This error quantifies how well the graph prompt approximates the desired graph data manipulation.
Empirical Maximum Error in Figure 1. Figure 1 in our paper illustrates the empirical maximum error observed in our experiments, corresponding to the theoretical upper bound discussed in Theorem 5. To approximate this upper bound: We perform multiple trials by repeating the experiment with different random removals of edges or nodes (i.e., generating different $G^{\prime}$) and different initializations of the graph prompt $P_{w}$. For each trial, we compute the error as described above. We record the maximum error observed across all trials, which serves as an empirical approximation of the theoretical upper bound.